OpenStudy (anonymous):

You role a standard number cube. find p grater then 4? 2/3 1/2 3/5 1/3

OpenStudy (anonymous):

Let \(X\) denote the number that shows up. Then \(P(X>4)=P(X=5,6)=P(X=5)+P(X=6)\).